OpenStudy (anonymous):

how do i prove that f(x)=sin(x)*cos(x) is one-to-one in -pi/3

OpenStudy (anonymous):

sorry can you give me the definition of one-to-one?

OpenStudy (anonymous):

hi. i mean that the following implication holds:\[if f(x _{1})= f(x _{2}), then x _{1}=x _{2}\]or its logical equivalent:\[if x _{1}\neq x _{2}, then f(x _{1})\neq f(x _{2})\]

OpenStudy (anonymous):

the one-to-ont aka(Bijection http://en.wikipedia.org/wiki/Bijection) can be proved given the function is cotinuous in that interval, plus the slope is always either positive or negative in that particular interval

OpenStudy (anonymous):

you could also use the trigonometric identity \[\sin (x)\times \cos(x)=1\div2 \times \sin(2\times x)\] and then compare \[\sin(2\times x1)=\sin(2\times x2)\] using the same ideas

OpenStudy (anonymous):

ok, thank you very much; i'll try both methods. by "one-to-one" i meant what wikipedia (and myself, in spanish (my first language)) calls an "injection"; a bijection is more than that, for it also requires the codomain to be the same as the image of f ("surjection"). i didn't know that the terms were ambiguous.